OpenStudy (anonymous):

Suppose we want to determine the (binomial) probability(p) of getting 5 heads in 14 flips of a 2-sided coin. Using the Binomial Probabilities Table in Appendix B of the text, what values of n, x and p would we use to look up this probability, and what would be the probability? I have probability (P) is .5 as it's a 50% chance of heads or tails. n=14 flips x=5 But I'm confused as to how to find. Please help?

OpenStudy (anonymous):

\[P(x=k)=\dbinom{n}{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

in your case since it is a fair coin, both \(p\) and \(1-p=\frac{1}{2}\)

OpenStudy (anonymous):

so the probability for having 5 flips out of 14 would be a 50% chance still?

OpenStudy (anonymous):

\(n=114,k=5\) you get \[\dbinom{14}{5}\left(\frac{1}{2}\right)^{14}\]

OpenStudy (anonymous):

i meant \(n=14\)

OpenStudy (anonymous):

\(\dbinom{14}{5}\) is the number of ways to choose 5 out of the 14 slots to put "heads" you probably have that button on your calculator, if not use this http://www.wolframalpha.com/input/?i=14+choose+5

OpenStudy (anonymous):

that will be your numerator, \(2^{14}\) is your denominator

OpenStudy (anonymous):

this is your exact answer http://www.wolframalpha.com/input/?i=%2814+choose+5%29%2F2^14

OpenStudy (anonymous):

thank you! I don't have a scientific calculator or none so I'm just a tad lost on this subject

OpenStudy (anonymous):

you are asked to look something up in a table, because apparently your book was written in like 1973 and the faculty has not seen fit to move in to the 21st century, either that or the administration refuses to adopt a new text book

OpenStudy (anonymous):

haha! isn't that the truth

OpenStudy (anonymous):

so to answer the question actually asked, \(n=14, x = 5,p=.5\)

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

yw